# Funnel Synthesis via LMI Copositivity Conditions for Nonlinear Systems

### v2

@@ Abstract
Funnel synthesis refers to a procedure for synthesizing a time-varying controlled invariant set and an associated control law around a nominal trajectory. The computation of the funnel involves solving a continuous-time differential equation or inequality, ensuring the invariance of the funnel. Previous approaches often compromise the invariance property of the funnel; for example, they may enforce the equation or the inequality only at discrete temporal nodes and do not have a formal guarantee of invariance at all times. This paper proposes a computational funnel synthesis method that can satisfy the invariance of the funnel without such compromises. We derive a finite number of linear matrix inequalities (LMIs) that imply the satifaction of a continuous-time differential linear matrix inequality guaranteeing the invariance of the funnel at all times from the initial to the final time. To this end, we utilize LMI conditions ensuring matrix copositivity, which then imply continuous-time invariance. The primary contribution of the paper is to prove that the resulting funnel is indeed invariant over a finite time horizon. We validate the proposed method via a three-dimensional trajectory planning and control problem with obstacle avoidance constraints.
@@ 1 Introduction
Funnels, also referred to as tubes in model predictive control (MPC) literature [1], represent a time-varying controlled invariant set around a nominal trajectory over a finite-time horizon. Funnel synthesis refers to a procedure not only computing the invariant funnel but also synthesizing an associated feedback control. The computed funnel with the associated controller can serve various purposes including robust motion planning and control [2] and trajectory planning with guaranteed feasibility [3].
@@
The proposed work synthesizes funnels based on uncertain linear time-varying (LTV) systems obtained via linearization of the original nonlinear dynamics around given nominal trajectories. Hence, the following papers on the analysis and synthesis of uncertain LTV systems is relevant to this paper: [8, 9, 10, 11]. The results in [9] provide a computational approach to analyze robust performance of uncertain LTV systems whose uncertainties are characterized by integral quadratic constraints (IQCs). This work is later extended in [8] to the synthesis of robust output-feedback controllers. The algorithm proposed in [8] alternately computes the variables describing the Lyapunov function for the closed-loop LTV system with a fixed feedback controller. Then it computes the feedback control gains while fixing the variables defining the underlying Lyapunov function until convergence. That is, they fix a set of solution variables while solving the rest, and then they solve for the solution variables that they fixed while the rest are now kept constant, and the process repeats in an iterative fashion. On the other hand, our proposed work characterizes nonlinearities using incremental quadratic constraints based on Lipschitz nonlinearities and constructs a single convex optimization problem whose solution provides both the funnel and feedback controller parameters.
@@ 1.1 Notation
The set notations $\mathbb{R}$, $\mathbb{R}_{+}$, $\mathbb{Z}$ and $\mathbb{R}^{n}$ are the sets of real, nonnegative real, integer, and the $n$-dimensional Euclidean spaces, respectively. Intervals are written by $\mathbb{Z}_{[a,b)}=\{z\in\mathbb{Z}:a\leq z<b\}$. The symmetric matrix $Q=Q^{\top}(\succeq)\succ 0$ implies $Q$ is a positive (semi) definite matrix. The set of positive (semi) definite matrices whose size is $n\times n$ are denoted by $\mathbb{S}_{++}^{n}(\mathbb{S}_{+}^{n})$. The identity matrix having $n\times n$ size is denoted by $I_{n}$. The subscript and the time argument will be omitted when it is clear from the context. The notation $\mathcal{L}_{2}[a,b]$ is a set of Lebesgue measurable functions $x(t)$ defined on an interval $[a,b]\subset\mathbb{R}$ such that $\left(\int_{a}^{b}x(t)^{\top}x(t)\text{d}t\right)^{1/2}<\infty.$ A property is said to hold almost everywhere, or for almost every $x\in X$ with some set $X$ if the set on which it fails is a Lebesgue measure zero set. We abbreviate the notations $A^{\top}PA$ and $\left[\begin{array}[]{cc}a&b^{\top}\\ b&c\end{array}\right]$ as $(\star)^{\top}PA$ and $\left[\begin{array}[]{cc}a&\star\\ b&c\end{array}\right]$, respectively. The Minkowski sum is denoted by $\oplus$.
@@ 2.1 Nonlinear systems
where $x(t)\in\mathbb{R}^{n_{x}}$ is the state, $u(t)\in\mathbb{R}^{n_{u}}$ is the input, $w(t)\in\mathbb{R}^{n_{w}}$ is the (exogenous) disturbance, and $t_{0},t_{f}$ are the initial and final time. The system dynamics $f:\mathbb{R}_{+}\times\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\times\mathbb{R}^{n_{w}}\rightarrow\mathbb{R}^{n_{x}}$ are assumed to be continuous in $t$ and continuously differentiable in $x,\,u$ and $w$. We assume that the input signal $u(\cdot)\in\mathcal{L}_{2}^{n_{u}}[t_{0},t_{f}]$ and the disturbance signal $w(\cdot)\in\mathcal{L}_{2}^{n_{w}}[t_{0},t_{f}]$ are piecewise continuous. We further assume that the disturbance signal $w(\cdot)$ is essentially bounded from above one, that is,
@@
where $\lVert w(\cdot)\rVert_{\infty}\coloneqq\text{esssup}_{t\in[t_{0},t_{f}]}\lVert w(t)\rVert_{2}$. We refer to a trajectory as a collection of the state, the input, and the disturbance signals, denoted together by $(x(\cdot),u(\cdot),w(\cdot))$. The nominal trajectory having the zero disturbance $(\bar{x}(\cdot),\bar{u}(\cdot),0)$ is assumed to be given and dynamically feasible, that is, $\dot{\bar{x}}=f(t,\bar{x},\bar{u},0)$ for all $t\in[t_{0},t_{f}]$. With the given nominal trajectory, we can rewrite the system (1) in the linear fractional form
@@
where $x_{1}$ and $x_{2}$ are $x$- and $y$-positions, $x_{3}$ are the yaw angle, $u_{1}$ is the velocity control, and $u_{2}$ is the angular velocity control. The disturbances $w_{1}$ and $w_{2}$ affect to the yaw angle estimation and the angular velocity control, and $c_{1},c_{2}\in\mathbb{R}_{+}$ are system parameters. The argument $q$ of $\phi$ can be chosen as $[x_{3},u_{1},w_{1}]$ with
@@
Since the only first two components in $f$ involve with the nonlinearity, the matrix $E_{o}$ is $\left[\begin{array}[]{ccc}1&0&0\\ 0&1&0\end{array}\right]^{\top}$. The time-varying matrices $A(t),\,B(t)$, and $F(t)$ are given as follows:
@@
where $\bar{x}=[\bar{x}_{1},\bar{x}_{2},\bar{x}_{3}]^{\top}$ adn $\bar{u}=[\bar{u}_{1},\bar{u}_{2}]^{\top}$ are the nominal state and input, respectively. The nonlinear function $\phi$ that is the remainder term in (3) is given by
@@ 2.2 Incremental dynamical system and its LPV approximation
for any compact set $\mathcal{Q}$. Here the term $\phi(t,q)-\phi(t,\bar{q})$, due to the nonlinearity of (3), is a state-, input-, and disturbance-dependent uncertainty characterized by the constraint (7).
Now we describe an LPV approximation of the LTV system (6). The underlying motivation of this approximation is that we can derive a finite number of LMIs that imply the invariance of the funnel for the LPV system, which is not tractable for the LTV system. We start by partitioning the time horizon $[t_{0},t_{f}]$ into $N\in\mathbb{R}_{+}$ uniform subintervals using time nodes defined as:
@@
On each subinterval $[t_{k},t_{k+1}]$, we approximate the time-varying system matrices $A(t)$, $B(t)$, and $F(t)$ using a first-order hold (FOH) approach, which linearly interpolates between their values at the endpoints of the interval. We define:
~$\tilde{\square}(t)$ $\coloneqq\sigma_{1}^{k}(t)\square_{k}+\sigma_{2}^{k}(t)\square_{k+1},\quad t\in[t_{k},t_{k+1}]$ (8a)
@@
where $\square_{k}=\square(t_{k})$ and the placeholder $\square$ corresponds to $A$, $B$, and $F$. By applying (8) across all $k$ in $\mathbb{Z}_{[0,N-1]}$, we obtain continuous piecewise linear approximations $(\tilde{A}(t),\tilde{B}(t),\tilde{F}(t))$ of the system matrices $(A(t),B(t),F(t))$ over the entire horizon $[t_{0},t_{f}]$.
@@
$\dot{\eta}$ $=\tilde{A}(t)\eta+\tilde{B}(t)\xi+\tilde{F}(t)w+\underbrace{\Delta(t)\left[\begin{array}[]{c}\eta\\ \xi\\ w\end{array}\right]}_{\coloneqq e(t)}+E_{o}\delta\phi,$ (9d)
$\Delta(t)$ $\coloneqq\left[\begin{array}[]{ccc}A(t)-\tilde{A}(t)&B(t)-\tilde{B}(t)&F(t)-\tilde{F}(t)\end{array}\right],$ (9f)
where $\delta\phi(t,\delta q)\coloneqq\phi(t,q)-\phi(t,\bar{q})$. Note that two representations (6) and (9) are equivalent by the definition of the error term $e(t)$ in (9d). Since $f$ is assumed to be continuously differentiable, the matrices $A(t)$, $B(t)$, and $F(t)$ are continuous, so bounded within the finite interval $[t_{0},t_{f}]$. Hence, there exists a positive constant $\beta(t)\in\mathbb{R}_{+}$ for each $t$ such that
$\lVert\Delta(t)\rVert_{2}\leq\beta(t),$ (10)
@@
where the matrix $E_{\Delta}\in\mathbb{R}^{n_{x}\times n_{\Delta}}$, the block-diagonal matrix $\Delta\in\mathbb{R}^{n_{\Delta}\times n_{\Delta}}$ of uncertainties, and the vector $q_{\Delta}\in\mathbb{R}^{n_{\Delta}}$ are given by
@@
with $n_{\Delta}=n_{q_{\eta}}+n_{q_{\xi}}+n_{q_{w}}$. These components are chosen such that
@@
$q_{\eta}=C_{\eta}\eta,\quad q_{\xi}=D_{\xi}\xi,\quad q_{w}G_{w}w.$
In this formulation, the uncertainties are captured using component-wise inequalities:
$|\Delta_{i}(t)|\leq\beta_{i}(t),\quad i=\mathbb{Z}_{[1,n_{\Delta}]},$ (11)
where $\beta_{i}(t)>0$ bounds the corresponding uncertainty $\Delta_{i}(t)$. Compared to the single norm-bounded inequality in (10), these component-wise bounds in (11) can represent the uncertainly more compactly by exploiting its structure. More details on uncertainty representation can be found in [17].
@@
$M(t)$ $=\left[\begin{array}[]{cc}N_{1}(t)^{-1}&0\\ 0&-N_{2}(t)^{-1}\end{array}\right],N_{1}(t)\coloneqq\left[\begin{array}[]{cc}\lambda_{\beta}(t)I_{n_{r}-n_{q}}&0\\ 0&\lambda_{\gamma}(t)I_{n_{q}}\end{array}\right],$ (14e)
$N_{2}(t)$ $\coloneqq\left[\begin{array}[]{cc}\beta(t)^{2}\lambda_{\beta}(t)I_{n_{\Delta}}&0\\ 0&\lambda_{\gamma}(t)\Gamma(t)^{2}\end{array}\right],$ (14h)
where $\Gamma(t)=\text{diag}(\gamma_{1}(t),\ldots,\gamma_{n_{\phi}}(t))$ and any positive real-valued functions $\lambda_{\beta}(t)>0$ and $\lambda_{\gamma}(t)>0$. This can be verified by deriving the following:
$\left[\begin{array}[]{c}r\\ p\end{array}\right]^{\top}M\left[\begin{array}[]{c}r\\ p\end{array}\right]$ $=\frac{1}{\lambda^{\beta}}\left(\beta^{2}(\eta^{\top}\eta+\xi^{\top}\xi+w^{\top}w)-e^{\top}e\right)$
$+\frac{1}{\lambda^{\gamma}}\left(\delta q^{\top}\Gamma(t)^{2}\delta q-\delta\phi^{\top}\delta\phi\right)\geq 0,$

### v3

@@ Abstract
Funnel synthesis refers to a procedure for synthesizing a time-varying controlled invariant set and an associated control law around a nominal trajectory. The computation of the funnel involves solving a continuous-time differential equation or inequality, ensuring the invariance of the funnel. Previous approaches often compromise the invariance property of the funnel; for example, they may enforce the equation or the inequality only at discrete temporal nodes and do not have a formal guarantee of invariance at all times. This paper proposes a computational funnel synthesis method that can satisfy the invariance of the funnel without such compromises. We derive a finite number of linear matrix inequalities (LMIs) that imply the satisfaction of a continuous-time differential linear matrix inequality guaranteeing the invariance of the funnel at all times from the initial to the final time. To this end, we utilize LMI conditions ensuring matrix copositivity, which then imply continuous-time invariance. The primary contribution of the paper is to prove that the resulting funnel is indeed invariant over a finite time horizon. We validate the proposed method via a three-dimensional trajectory planning and control problem with obstacle avoidance constraints, and a six-degree-of-freedom powered descent guidance.
@@ 1 Introduction
Funnels, also referred to as tubes in model predictive control (MPC) literature [1, 2], represent a time-varying controlled invariant set around a nominal trajectory over a finite-time horizon. Funnel synthesis refers to a procedure not only computing the invariant funnel but also synthesizing an associated feedback control. The computed funnel with the associated controller can serve various purposes including robust motion planning and control [3] and trajectory planning with guaranteed feasibility [4].
@@
The proposed work synthesizes funnels based on uncertain linear time-varying (LTV) systems obtained via linearization of the original nonlinear dynamics around given nominal trajectories. Hence, the following papers on the analysis and synthesis of uncertain LTV systems are relevant to this paper: [9, 10, 11, 12]. The results in [10] provide a computational approach to analyze robust performance of uncertain LTV systems whose uncertainties are characterized by integral quadratic constraints (IQCs). This work is later extended in [9] to the synthesis of robust output-feedback controllers. The algorithm proposed in [9] alternately computes the variables describing the Lyapunov function for the closed-loop LTV system with a fixed feedback controller. Then it computes the feedback control gains while fixing the variables defining the underlying Lyapunov function until convergence. That is, they fix a set of solution variables while solving the rest, and then they solve for the solution variables that they fixed while the rest are now kept constant, and the process repeats in an iterative fashion. On the other hand, our proposed work characterizes nonlinearities using incremental quadratic constraints based on Lipschitz nonlinearities and constructs a single convex optimization problem whose solution provides both the funnel and feedback controller parameters.
@@ 1.1 Notation
The set notations $\mathbb{R}$, $\mathbb{R}_{+}$, $\mathbb{Z}$ and $\mathbb{R}^{n}$ are the sets of real, nonnegative real, integer, and the $n$-dimensional Euclidean spaces, respectively. Intervals are written by $\mathbb{Z}_{[a,b)}=\{z\in\mathbb{Z}:a\leq z<b\}$. The symmetric matrix $Q=Q^{\top}(\succeq)\succ 0$ implies $Q$ is a positive (semi) definite PD (PSD) matrix. The set of positive (semi) definite matrices whose size is $n\times n$ are denoted by $\mathbb{S}_{++}^{n}(\mathbb{S}_{+}^{n})$. The identity matrix having $n\times n$ size is denoted by $I_{n}$. The subscript and the time argument will be omitted when it is clear from the context. The notation $\mathcal{L}_{2}[a,b]$ is a set of Lebesgue measurable functions $x(t)$ defined on an interval $a,b]\subset\mathbb{R}$ such that $\left(\int_{a}^{b}x(t)^{\top}x(t)\text{d}t\right)^{1/2}<\infty.$ A property is said to hold almost everywhere, or for almost every $x\in X$ with some set $X$ if the set on which it fails is a Lebesgue measure zero set. We abbreviate the notations $A^{\top}PA$ and $\left[\begin{array}[]{cc}a&b^{\top}\\ b&c\end{array}\right]$ as $(\star)^{\top}PA$ and $\left[\begin{array}[]{cc}a&\star\\ b&c\end{array}\right]$, respectively. The Minkowski sum is denoted by $\oplus$. The stacking of vectors $x$ and $y$ is denoted by $(x,y)=[x^{\top}y^{\top}]^{\top}$. The operation $\text{diag}(\cdot)$ outputs a diagonal matrix formed from a given vector.
@@ 2.1 Nonlinear systems
where $x(t)\in\mathbb{R}^{n_{x}}$ is the state, $u(t)\in\mathbb{R}^{n_{u}}$ is the input, $w(t)\in\mathbb{R}^{n_{w}}$ is the (exogenous) disturbance, and $t_{0},t_{f}$ are the initial and final time. The system dynamics $f:\mathbb{R}_{+}\times\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\times\mathbb{R}^{n_{w}}\rightarrow\mathbb{R}^{n_{x}}$ are assumed to be continuous in $t$ and continuously differentiable in x,ux,\,u and $w$. We assume that the input signal $u(\cdot)\in\mathcal{L}_{2}^{n_{u}}[t_{0},t_{f}]$ and the disturbance signal $w(\cdot)\in\mathcal{L}_{2}^{n_{w}}[t_{0},t_{f}]$ are piecewise continuous. We further assume that the disturbance signal $w(\cdot)$ is essentially bounded from above by one, that is,
@@
where $\lVert w(\cdot)\rVert_{\infty}\coloneqq\text{esssup}_{t\in[t_{0},t_{f}]}\lVert w(t)\rVert_{2}$. Notice that the choice of the upper bound (i.e., one) is not restrictive, since if $w(\cdot)$ is bounded above by some constant $w_{max}\in\mathbb{R}_{+}$, we can redefine $w\leftarrow w/w_{max}$ so that the bound becomes one. We refer to a trajectory as a collection of the state, the input, and the disturbance signals, denoted together by $(x(\cdot),u(\cdot),w(\cdot))$. The nominal trajectory having the zero disturbance $(\bar{x}(\cdot),\bar{u}(\cdot),0)$ is assumed to be given and dynamically feasible, that is, $\dot{\bar{x}}=f(t,\bar{x},\bar{u},0)$ for all $t\in[t_{0},t_{f}]$. With the given nominal trajectory, we can rewrite the system (1) in the linear fractional form
@@
where $x_{1}$ and $x_{2}$ are $x$- and $y$-positions, $x_{3}$ are the yaw angle, $u_{1}$ is the velocity control, and $u_{2}$ is the angular velocity control. The disturbances $w_{1}$ and $w_{2}$ affect to the yaw angle estimation and the angular velocity control, and $c_{1},c_{2}\in\mathbb{R}_{+}$ are system parameters. The argument $q$ of $\phi$ can be chosen as $x_{3},u_{1},w_{1}]$ with
@@
Since the only first two components in $f$ involve with the nonlinearity, the matrix $E_{o}$ is $\left[\begin{array}[]{ccc}1&0&0\\ 0&1&0\end{array}\right]^{\top}$. The time-varying matrices A(t),B(t)A(t),\,B(t), and $F(t)$ are given as follows:
@@
where $\bar{x}=[\bar{x}_{1},\bar{x}_{2},\bar{x}_{3}]^{\top}$ and $\bar{u}=[\bar{u}_{1},\bar{u}_{2}]^{\top}$ are the nominal state and input, respectively. The nonlinear function $\phi$ that is the remainder term in (3) is given by
@@ 2.2 Incremental dynamical system and its LPV approximation
for any compact set $\mathcal{Q}\subseteq\mathbb{R}^{n_{q}}$. Here the term $\phi(t,q)-\phi(t,\bar{q})$, due to the nonlinearity of (3), is a state-, input-, and disturbance-dependent uncertainty characterized by the constraint (7). This type of LTV system in incremental form with Lipschitz nonlinearity has been studied in the context of funnel computation in several existing works. See, for example, [4, 6] for related formulations and applications.
Now we describe an LPV approximation of the LTV system (6). The underlying motivation of this approximation is that we can derive a finite number of LMIs that imply the invariance of the funnel for the LPV system, which is not tractable for the LTV system. We start by partitioning the time horizon $t_{0},t_{f}]$ into $N\in\mathbb{R}_{+}$ uniform subintervals using time nodes defined as:
@@
On each subinterval $t_{k},t_{k+1}]$, we approximate the time-varying system matrices $A(t)$, $B(t)$, and $F(t)$ using a first-order hold (FOH) approach, which linearly interpolates between their values at the endpoints of the interval. This results in convex combinations of the matrices at $t_{k}$ and $t_{k+1}$. We define:
$\tilde{\square}(t)$ $\coloneqq\sigma_{1}^{k}(t)\square_{k}+\sigma_{2}^{k}(t)\square_{k+1},\quad t\in[t_{k},t_{k+1}]$ (8a)
@@
where $\square_{k}=\square(t_{k})$ and the placeholder $\square$ corresponds to $A$, $B$, and $F$. By applying (8) across all $k$ in $\mathbb{Z}_{[0,N-1]}$, we obtain continuous piecewise linear approximations $(\tilde{A}(t),\tilde{B}(t),\tilde{F}(t))$ of the system matrices $(A(t),B(t),F(t))$ over the entire horizon $t_{0},t_{f}]$.
It is worth noting that the system matrices $(A(t),B(t),F(t))$ are not necessarily piecewise linear, and therefore the approximation error is generally nonzero. However, it is common in the literature (e.g., [4, 18]) to assume piecewise linearity to simplify funnel computation. In contrast, this paper does not assume zero approximation error; instead, the error is modeled explicitly as a state-, input-, and disturbance-dependent uncertainty.
@@
$\dot{\eta}$ $=\tilde{A}(t)\eta+\tilde{B}(t)\xi+\tilde{F}(t)w+\underbrace{\Delta_{e}(t)\left[\begin{array}[]{c}\eta\\ \xi\\ w\end{array}\right]}_{\coloneqq e(t)}+E_{o}\delta\phi,$ (9d)
$\Delta_{e}(t)$ $\coloneqq\left[\begin{array}[]{ccc}A(t)-\tilde{A}(t)&B(t)-\tilde{B}(t)&F(t)-\tilde{F}(t)\end{array}\right],$ (9f)
where $\delta\phi(t,\delta q)\coloneqq\phi(t,q)-\phi(t,\bar{q})$. Note that two representations (6) and (9) are equivalent by the definition of the error term $e(t)$ in (9d). Since $f$ is assumed to be continuously differentiable, the matrices $A(t)$, $B(t)$, and $F(t)$ are continuous, so bounded within the finite interval $t_{0},t_{f}]$. Hence, there exists a positive constant $\beta_{e}(t)\in\mathbb{R}_{+}$ for each $t$ such that
$\lVert\Delta_{e}(t)\rVert_{2}\leq\beta_{e}(t),$ (10)
@@
where the matrix $E_{\Delta}\in\mathbb{R}^{n_{x}\times n_{\Delta}}$, the block-diagonal matrix $\Delta\in\mathbb{R}^{n_{\Delta}\times n_{q_{\Delta}}}$ of uncertainties, and the vector $q_{\Delta}\in\mathbb{R}^{n_{q_{\Delta}}}$ are given by
@@
with $n_{\Delta}\in\mathbb{R}_{+}$ and $n_{q_{\Delta}}=n_{q_{\eta}}+n_{q_{\xi}}+n_{q_{w}}$. These components are chosen such that
@@
$q_{\eta}=C_{\eta}\eta,\quad q_{\xi}=D_{\xi}\xi,\quad q_{w}=G_{w}w.$
In this formulation, the uncertainties are captured using the following inequalities:
$\lVert\Delta_{\square}(t)\rVert_{2}\leq\beta_{\square}(t),$ (11)
where $\beta_{\square}(t)>0$ bounds the corresponding uncertainty $\Delta_{\square}(t)$, with the placeholder $\square$ representing the subscripts $\eta$, $\xi$, and $w$. Compared to the single norm-bounded inequality in (10), these bounds in (11) can represent the uncertainly more compactly by exploiting its structure. More details on the structured uncertainty representation can be found in [19, 6.2.1], [20], [21].
@@
$M(t)$ $=\left[\begin{array}[]{cc}N_{1}(t)^{-1}&0\\ 0&-N_{2}(t)^{-1}\end{array}\right],N_{1}(t)\coloneqq\left[\begin{array}[]{cc}N_{1}^{\beta}(t)&0\\ 0&\frac{1}{n_{\phi}}\lambda_{\gamma}(t)I_{n_{q}}\end{array}\right],$ (14e)
$N_{2}(t)$ $\coloneqq\left[\begin{array}[]{cc}N_{2}^{\beta}(t)&0\\ 0&\lambda_{\gamma}(t)\Gamma(t)^{2}\end{array}\right],N_{1}^{\beta}(t)\coloneqq\text{diag}(\lambda_{\beta_{\eta}}(t)I,\lambda_{\beta_{\xi}}(t)I,\lambda_{\beta_{w}}(t)I),$ (14h)
$N_{2}^{\beta}(t)$ $\coloneqq\text{diag}(\lambda_{\beta_{\eta}}(t)\beta_{\eta}^{2}(t)I,\lambda_{\beta_{\xi}}(t)\beta_{\xi}^{2}(t)I,\lambda_{\beta_{w}}(t)\beta_{w}^{2}(t)I),$ (14i)
where $\Gamma(t)=\text{diag}(\gamma_{1}(t),\ldots,\gamma_{n_{\phi}}(t))$ and any positive real-valued functions $\lambda_{\beta_{\square}}(t)>0$ with the placeholder $\square$ representing $\eta$, $\xi$, and $w$, and $\lambda_{\gamma}(t)>0$. This can be verified by deriving the following:
$\left[\begin{array}[]{c}r\\ p\end{array}\right]^{\top}M\left[\begin{array}[]{c}r\\ p\end{array}\right]$ $=\sum_{\square=\{\eta,\xi,w\}}\frac{1}{\lambda_{\beta_{\square}}}(q_{\square}^{\top}q_{\square}-\frac{1}{\beta_{\square}^{2}}q_{\square}^{\top}\Delta_{\square}^{\top}\Delta_{\square}q_{\square})$
$+\frac{1}{\lambda_{\gamma}}\left(n_{\phi}\delta q^{\top}\delta q-\delta\phi^{\top}\Gamma(t)^{-2}\delta\phi\right)\geq 0,$